What is the func. of reagent SeO2 ? Pls giv an eg.

Grade:12

3 Answers

nikhil arora
12 years ago

its an oxidising agent used in many organic reactions.

e.g. 

*  used for oxidation at allylic carbon.

    CH3CH2CH=CH2 +SeO2 -------> CH3CH(OH)CH=CH2

Shyama P
12 years ago

Is it a mild oxi. agent?

Karun
5 years ago
SeO2 also oxidises -CH3 or       -CH2 in the alpha position (with respect to the -(C=O) present) to  -CHO and -(C=O)- respectively. 
eg:- CH3-(C=O)-H --> O=CH-CH=O (Glyoxal).